A Mars satellite moving in an orbit of radius 9.4 × 103 km takes 27540 s to complete one revolution. Calculate the mass of Mars.

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol. T = 2π √r^3/GM ⇒ 27540 = 2 * 3.14 √(9.4 * 10^3 * 10^3)^3/6.67 * 10^-11 * M Or (27540)^2 = (6.28)^2 (9.4 * 10^6)^2/6.67 * 10^-11 *M Or M = (6.28)^2 * (9.4)^3 * 10^18/6.67 * 10^-11 * (27540)^2 = 6.5 * 10^23 kg.